Calculate the mass % of each component in the mixture based on the total mass recovered?( Mass of sand + beaker= 117.8120, mass of salt +1st beaker= 116.7434, mass of salt + 2nd beaker= 117.4350)
1
Expert's answer
2019-10-14T14:30:41-0400

Here,

Let the mass of sand be "x"

Mass of 1st beaker be "y"

Mass of 2nd beaker be "z"

Then

Mass of sand + beakers = "x+y+z=117.8120\ \ \ \ .......(1)"

Similarly,

"x+y=116.7434 \ \ \ \ \ .....(2)\\ x+z=117.4350\ \ \ \ \ .........(3)"

Putting the value of (3) in (1)

We get,

"y=0.3770"

Putting the value of y in (2)

We get,

"x=116.3664"

Putting the value of x in (3)

We get ,

"z=1.0686"

Total mass = "117.8120"

Mass percentage of sand

"=\frac{116.3664×100}{117.8120}=98.773\ ℅"

Mass percentage of beaker 1 =

"=\frac{0.3770×100}{117.8120}=0.32\ ℅"

Mass percentage of beaker 2

"=\frac{1.0686×100}{117.8120}=0.907\ ℅"